OpenStudy (jojokiw3):

A rat was thrown at an angle of 37° with a speed of 10 m/s. How long is it in the air?

OpenStudy (agent0smith):

Just...why was a rat thrown? What a bizarre scenario. If this is a physics question, you can use equations of motion to solve it. First find the y component of velocity, which is 10sin37. Then use \[\Large v_f = v_0 + at\] since you can solve for when the velocity is zero. Don't forget to double it, since when v reaches zero, it's only halfway through it's flight.

OpenStudy (jojokiw3):

t = 1.2 s.

OpenStudy (agent0smith):

Show your work.

OpenStudy (jojokiw3):

\[0 = 0 + v_{z0}t - \frac{ 1 }{ 2 }g t^2\] \[t = \frac{ 2v_{z0} }{ g } = 2(10m/s) \sin 37°/ (9.8 m/s^2)\]

OpenStudy (agent0smith):

Looks right.

OpenStudy (agent0smith):

Seriously who the hell throws a rat? Was it alive?

OpenStudy (jojokiw3):

:D YES.

OpenStudy (agent0smith):

Maybe it was thrown to kill it then.

OpenStudy (jojokiw3):

Not only does it rid the world of a possible vermin, it also teaches physics. I'd say the rat lived a valuable life.

OpenStudy (agent0smith):

Agreed.

OpenStudy (jojokiw3):

Thanks for answering the question! :D

OpenStudy (agent0smith):

You're welcome. The main reason I even clicked on it was because I thought it said "A rat was thrown at a wall at an angle of..."